# ScribblePrompt: Fast and Flexible Interactive Segmentation for Any Medical Image

### v1

@@ 3.3 Data
We build on MegaMedical [20] to construct a collection of 77 open-access medical imaging datasets for training and evaluation. The collection includes a diverse array of biomedical domains, such as eyes [48, 69, 83, 102, 122], thorax [112, 116, 118, 114, 103], spine vertebrae [148, 78, 114, 134], cells [149, 77, 21, 36, 22, 35], skin [26], abdominal [15, 16, 45, 54, 57, 65, 66, 68, 74, 79, 103, 118, 84, 121, 107], neck [62, 101, 95, 99], brain [10, 37, 46, 63, 64, 86, 87, 89, 118, 4], bones [114, 42, 134], teeth [1, 51] and lesions [5, 146, 147, 118].
We define a 2D segmentation task as a combination of dataset, axis (for 3D modalities), and label. For datasets with multiple segmentation labels, we consider each label separately as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice for each volume.
Task Diversity. During training we sample image, segmentation pairs hierarchically i.e., by dataset and modality, axis and then label, to balance training on datasets of different sizes. To increase the diversity of segmentation tasks, we apply data augmentation to both the input image and sampled segmentation prior to simulating the user interactions.
Synthetic Labels. To limit overfitting to specific segmentation tasks, we introduce synthetic labels. During training, for a given sample $(x^{t},y^{t})$, with probability $p_{synth}$ we replace $y^{t}$ with a synthetic label $y^{\prime}$. Given image $x^{t}$ we generate a synthetic label $y^{\prime}$ by applying a superpixel algorithm [33] with scale parameter $\lambda\sim U[1,\lambda_{max}]$ to partition the image into a multi-label mask of $k$ superpixels $z\in\{1,\dots,k\}^{n\times n}$ and then randomly selected a superpixel $y^{\prime}:=\mathbbm{1}(z=k)$.
@@ 3.4 Network
We demonstrate ScribblePrompt using an efficient fully-convolutional architecture similar to a Unet [108], ScribblePrompt-Unet, and also a vision transformer-based architecture, the Segment Anything Model (SAM) [61], ScribblePrompt-SAM. We provide code and weights for both models at scribbleprompt.csail.mit.edu.
@@
ScribblePrompt-SAM. For ScribblePrompt-SAM we fine-tune the smallest SAM model (ViT-b). We use ViT-b because we observed the smallest backbone (ViT-b) had similar performance to the larger backbones on the validation set. We freeze all weights except for the decoder.
The SAM architecture takes bounding box and click inputs as lists of ${(x,y)}$ coordinates, and the low-resolution logits of the previous prediction $y_{i-1}$. To adapt SAM for scribbles, we represent positive/negative scribbles as a set of positive/negative clicks on every pixel of the scribbles.
Loss. We minimize eq. (1) where $\mathcal{L}_{Seg}$ is a linear combination of soft Dice Loss [32] and Focal Loss [72]. In preliminary experiments (Appendix B.2), training with Soft Dice Loss [32] alone or a linear combination of Soft Dice Loss and Binary-Cross Entropy Loss resulted in slightly lower Dice scores.
@@ 4 Experimental Setup
We compare ScribblePrompt-Unet and ScribblePrompt-SAM to previous interactive segmentation methods with both a user study with experienced annotators, and with manual and simulated interactions.
@@
Data. We use 65 datasets during training. We partition another nine datasets, ACDC [14], BUID [5], BTCV Cervix [66], DRIVE [122], HipXRay [42], PanDental [1], SCD [103], SpineWeb [148], and WBC [149], into a validation set (used in model selection, but not training) and a test set (used only for final evaluation). Each dataset was split into 60% train, 20% validation, and 20% test by subject. Three additional datasets, TotalSegmentator [134], SCR [125], and COBRE [4], were used only for final evaluation. We selected these datasets to cover a variety of modalities (MRI, CT, Ultrasound, OCT, Microscopy) and anatomical regions (Brain, Teeth, Bones, Organs, Muscles, Heart, Thorax, Cells), including both healthy anatomy and lesions. We report results on twelve sets of data covering 608 tasks and eight modalities.
We resized all images to 128x128 and re-scaled intensities to [0,1]. For methods using the SAM architecture, we convert the images to RGB and apply the pixel normalization scheme in [61]. We train with $p_{synth}=0.5$ and apply the same data augmentations as [20]. We provide more details on the data in Appendix F.
@@
Training. We simulate five iterations of interactions per example and set the maximum number of interactions per iteration to three. We set the minimum number of initial negative prompts to zero, and the minimum initial positive and correction prompts to one. During the first iteration, we sample the combination of prompt types (clicks, scribbles, bounding boxes) and then sample the number of positive and negative interactions for each type. In each subsequent iteration, we simulate either scribble corrections or click corrections with equal probability. We selected these values based on what we believe is a reasonable number of steps and interactions a user might perform.
We use the Adam optimizer [59] and train with a learning rate of $0.0001$ until convergence. We use a batch size of 8 for ScribblePrompt-Unet. For ScribblePrompt-SAM we use a batch size of 1, because of memory constraints.
Figure 3: Example predictions. Left to right: SAM, SAM-Med2d, MIDeepSeg, ScribblePrompt-Unet, and ScribblePrompt-SAM. Top: predictions after one step of manual scribbles. Bottom: predictions after five steps of simulated interactions (one center click or bounding box, followed by one center correction click per step). Representative examples were selected to show the diversity of tasks in our evaluation.
4.1 Baselines
We compare to the following existing methods for iterative interactive segmentation of medical images.
@@ 4.1 Baselines
SAM [61]: We evaluate the smallest (ViT-b) and largest (ViT-h) versions of the Segment Anything Model (SAM), trained on a large natural image dataset. SAM takes bounding boxes and clicks as input.
@@
SAM-Med2d [25]: SAM-Med2d is a SAM model with ViT-b backbone and additional adapter layers in the image encoder. SAM-Med2d was fine-tuned on a collection of public and private datasets containing 4.6M medical images and 19.7M segmentations [^1] SAM-Med2d used three of our evaluation datasets (ACDC, BTCV and TotalSegmentator) as training datasets [140]., using bounding boxes and iterative clicks as input. Following [25], we evaluate SAM-Med2d both with and without adapter layers.
@@
MIDeepSeg [80]: MIDeepSeg is an interactive segmentation framework designed to generalize to unseen tasks on medical images. MIDeepSeg takes interior margin points (positive clicks) as initial inputs. It crops the image based on those points and uses a CNN to make an initial prediction. The prediction can be refined with additional clicks using conditional random fields. We evaluate the pre-trained model, which was developed on Placenta T2 MRI.
5 Experiments
@@
We evaluate all methods using both manual and simulated interactions. Simulated interactions enable us to test on a large volume of images and tasks. However, simulations do not always match user behavior. We address this by (1) collecting a small dataset of manual scribbles for validation and (2) conducting an interactive user study.
@@ 5.1 Manual Scribbles
Setup. To validate the proposed scribble simulation procedure, we collected a dataset of manual scribbles from three annotators for seven segmentation tasks with a total of 93 examples. The annotators were shown five training examples with the ground truth segmentation per task and instructed to draw positive and negative scribbles to indicate the region of interest on new images.
In this evaluation, we make a prediction given a set of positive and negative scribbles, simulating a non-iterative scenario where the annotator draws several scribbles before running inference. We evaluate accuracy using Dice score [32]. MIDeepSeg’s initial prediction network only takes positive inputs; we report results on both MIDeepseg’s initial prediction from the positive scribbles, and after applying MIDeepSeg’s refinement procedure using the negative scribbles as if they were corrections.
Results. ScribblePrompt-SAM and ScribblePrompt-Unet produce the most accurate segmentations in a single step from the manual scribbles, with an average Dice of 0.87 and 0.81, respectively (Table 1). Scribbles are particularly helpful for large segmentations covering areas with varying intensity, such as the jaw in dental X-Rays (Figure 3).
SAM and SAM-Med2d do not generalize well to scribble inputs (which they were not trained for). The initial predictions from MIDeepSeg’s network are poor (0.68 Dice), but improve to 0.81 Dice after applying the refinement procedure, suggesting that for some examples in our manual scribble dataset non-DL methods may be sufficient. We evaluate the generalizability of all methods on more tasks through experiments with simulated interactions.
Table 1: Evaluation on manual scribbles. Mean Dice of segmentations ($n=93$) predicted from manual scribbles on examples from validation datasets. The ScribblePrompt models, which were trained with simulated scribbles (as well as clicks and boxes), have the highest Dice from a single prediction.
Model Dice (mean $\pm$ ci)
SAM (ViT-b) $0.40\pm 0.05$
SAM (ViT-h) $0.57\pm 0.05$
SAM-Med2d w.o. adapter $0.52\pm 0.05$
SAM-Med2d w/ adapter $0.30\pm 0.06$
MIDeepSeg $0.68\pm 0.04$
MIDeepSeg (with refinement) $0.81\pm 0.03$
ScribblePrompt-Unet $0.81\pm 0.03$
ScribblePrompt-SAM $\bm{0.87\pm 0.02}$
@@ 5.2 Simulated Interactions
Setup. We use the scribble and click simulator to evaluate the performance of ScribblePrompt and baselines with iterative scribbles across the 12 evaluation sets.
@@
We limit each scribble to cover a maximum of 128 pixels. For centerline and contour scribbles, a scribble may contain multiple disconnected components.
@@
  • Warm Start: Three random positive clicks and three random negative clicks to start, followed by one (positive or negative) correction click per step in the center of the largest component of the error region.
Since MIDeepSeg crops the input image based on the input clicks (or scribbles), it cannot make accurate predictions with only a few clicks, so we exclude it from simulations that start with a single click. For each example and interaction procedure, we simulated series of iterative interactions (10 steps for scribble-focused procedures and 20 steps for click-focused procedures) with five random seeds.
Results. As shown in Figure 4, both versions of ScribblePrompt outperform baseline methods for all simulated interaction procedures at all numbers of interactions. We show qualitative examples with simulated interactions in Figure 1 and Figure 3. We show additional results with simulated bounding boxes in Appendix D.2.
Discussion. Comparing the simulated scribble and click results (Figure 4) highlights the efficiency of scribble-based interactions. Both ScribblePrompt models reach an average Dice above $0.8$ in two scribble steps. It takes five steps of carefully-placed center clicks or seven steps of random clicks to reach the same average Dice with ScribblePrompt-SAM (Figure 18). ScribblePrompt-SAM is more robust to random clicks than ScribblePrompt-Unet, perhaps because of its pre-training with random clicks on a large number of natural images.
On average, the MIDeepSeg first prediction is poor, because the network only takes positive inputs and was trained with interior margin point clicks on a single modality. Although SAM-Med2d was trained on a large corpus of medical imaging datasets, it does not perform as well as SAM on our evaluation sets. For some tasks, such as organ segmentation, SAM-Med2d, is very accurate. However, in cases where the initial prediction is poor (e.g., retinal veins, teeth in dental X-Rays), it is less accurate than SAM or MIDeepSeg when revising its predictions based on (simulated) corrections. Detailed results by evaluation set are provided in Appendix D.3.
Figure 4: Results with simulated clicks and scribbles. We evaluate methods by simulating interactions following three scribble-focused protocols and three click-focused protocols. We show the other results, which have similar trends, in Appendix D.3. We measure Dice averaged across twelve evaluation sets, weighting each dataset equally. Shaded regions show 95% CI.
6 User Study
We conducted a user study comparing the ease of use of ScribblePrompt-Unet to SAM (ViT-b). We selected ScribblePrompt-Unet for the user study because it had similar performance and lower latency compared to ScribblePrompt-SAM. We selected SAM (ViT-b) as the baseline because it had the highest Dice in our experiments with simulated clicks and lower latency than the next closest baseline, SAM (ViT-h).
Setup. Study participants ($n=16$) were neuroimaging researchers at an academic hospital. They used each model to segment a series of nine images from nine different evaluation sets. For each segmentation task, the participants were shown the target segmentation and were asked to prompt the model until the predicted segmentation closely matched the target or they could no longer improve the prediction. We provided participants with the target segmentation to disentangle the cognitive process of identifying the region of interest from prompting the model to achieve the desired segmentation. We provide additional details and visualizations in Appendix C.
Table 2: Quantitative results of user study. We report mean Dice score and standard deviation, mean time and standard deviation (in seconds) per segmentation, and median (max) number of times the participants refreshed the prediction (per segmentation task).
SAM (ViT-b) ScribblePrompt

### v2

@@ 3.3 Data
We build on large dataset gathering efforts like MegaMedical [20, 109] to compile a collection of 77 open-access biomedical imaging datasets for training and evaluation, covering over 54k scans, 16 image types, and 711 labels. The collection includes a diverse array of biomedical domains, such as eyes [49, 72, 85, 104, 125], thorax [115, 119, 121, 117, 105], spine [150, 80, 117, 137], cells [151, 79, 21, 37, 22, 36], skin [25], abdominal [15, 16, 46, 56, 59, 67, 68, 71, 76, 81, 105, 121, 86, 124, 110], neck [64, 103, 97, 101], brain [10, 38, 47, 65, 66, 88, 89, 91, 121, 4], bones [117, 43, 137], teeth [1, 53] and lesions [5, 148, 149, 121].
We define a 2D segmentation task as a combination of dataset, axis (for 3D modalities), and label. For datasets with multiple segmentation labels, we consider each label separately as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice from each volume. We provide more details on the data in Appendix C.
Task Diversity. During training we sample image, segmentation pairs hierarchically – by dataset and modality, axis, and then label – to balance training on datasets of different sizes. To increase the diversity of segmentation tasks, we apply data augmentation [20] to both the input image and sampled segmentation prior to simulating the user interactions.
Figure 4: Task sampling and augmentation. During training, we sample an image and segmentation pair $(x_{0},y_{0})$. With probability $p_{synth}$, we replace $y_{0}$ with a synthetic label $y_{synth}$. We generate $y_{synth}$ by applying a superpixel algorithm to the image $x_{0}$ to generate a map $z$ of potential synthetic labels (superpixels), and then sampling one label. Finally, we apply random data augmentations to get $(x^{t},y^{t})$.
Synthetic Labels. To limit overfitting to specific segmentation tasks, we introduce a mechanism to generate synthetic labels (Fig. 4). During training, for a given sample $(x_{0},y_{0})$, with probability $p_{synth}$, we replace $y$ with a synthetic label $y_{synth}$. Given an image $x_{0}$, we generate a synthetic label $y_{synth}$ by applying a superpixel algorithm [33] with scale parameter $\lambda\sim U[1,\lambda_{max}]$ to partition the image into a multi-label mask of $k$ superpixels $z\in\{1,\dots,k\}^{n\times n}$, and then randomly select a superpixel $y_{synth}=\mathbbm{1}(z=c)$. We show examples in Fig. 14 and conduct experiments varying $p_{synth}$.
@@ 3.4 Network
Motivated by producing a practical tool, we primarily demonstrate ScribblePrompt using an efficient fully-convolutional architecture similar to a UNet [111]. We also demonstrate ScribblePrompt using a vision transformer-based architecture, SAM ViT-b [63]. We provide code and weights for both models at scribbleprompt.csail.mit.edu.
@@
We encode each prompt type in an input channel. The input to ScribblePrompt-UNet has size $5\times h\times w$ consisting of the input image $x^{t}$, bounding box encoding, positive click/scribble encoding, negative click/scribble encoding, and the logits of the previous prediction $\hat{y}^{t}_{i-1}$. We encode bounding boxes in a binary mask that is 1 inside the box(es) and 0 everywhere else. We encode positive and negative clicks using binary masks, and combine them with the positive and negative scribble masks.
ScribblePrompt-SAM. We take the smallest SAM model (ViT-b) and fine-tune the decoder. SAM takes bounding box and click inputs as lists of ${(x,y)}$ coordinates, and the low-resolution logits of the previous prediction $\hat{y}_{i-1}$. To adapt SAM for scribbles, we consider each positive/negative scribbled pixel as a positive/negative click.
Loss. We minimize eq. (1) where $\mathcal{L}_{Seg}$ is a linear combination of soft Dice Loss [32] and Focal Loss [74]. In preliminary experiments (Sec. B.2), training with Soft Dice Loss [32] alone or a linear combination of Soft Dice Loss and Binary-Cross Entropy Loss resulted in slightly lower Dice scores.
@@ 4 Experimental Setup
We compare ScribblePrompt-UNet and ScribblePrompt-SAM to previous methods through experiments with manual scribbles, simulated interactions, and a user study with experienced annotators. Lastly, we report on inference runtime and ablation experiments.
@@
Data. We use 65 (out of 77) datasets during training. We partition another nine datasets, ACDC [14], BUID [5], BTCV Cervix [68], DRIVE [125], HipXRay [43], PanDental [1], SCD [105], SpineWeb [150], and WBC [151], each into validation (used in model selection, but not training) and test (used only for final evaluation). Three additional datasets, TotalSegmentator [137], SCR [128], and COBRE [4, 34, 29], were used only for final evaluation.
We report final results on 12 sets of data (the test splits of the 9+3 evaluation datasets). The evaluation datasets were not seen during training, and cover 608 tasks and 8 modalities, including unseen image types and unseen labels. We selected these 12 evaluation datasets to cover a variety of modalities (MRI, CT, ultrasound, fundus photography, microscopy) and anatomical regions of interest (brain, teeth, bones, abdominal organs, muscles, heart, thorax, cells), including both healthy anatomy and lesions.
@@
Training. During training, we simulated five steps of interactive segmentation per example and set the maximum number of interactions per step to three. We set the minimum number of initial negative prompts to zero, and the minimum initial positive and correction prompts to one. During the first step, we sample the combination of prompt types (clicks, scribbles, bounding boxes) and then sample the number of positive and negative interactions for each type. In each subsequent steps, we simulate either scribble corrections or click corrections with equal probability. We selected these values based on what we believe to be reasonable interactions for a user to perform.
Baselines. We compare to existing generalist methods for interactive segmentation, with a focus on methods developed for biomedical imaging.
@@
SAM [63]: We evaluate the smallest (ViT-b) and largest (ViT-h) versions of the Segment Anything Model (SAM) trained on natural images. SAM takes bounding boxes, clicks, and the logits of the previous prediction as input.
@@
SAM-Med2D [24]: SAM-Med2D is a SAM ViT-b model with additional adapter layers in the image encoder. SAM-Med2D was fine-tuned, using bounding boxes and iterative clicks as input, on a collection of biomedical imaging datasets containing 4.6M images and 19.7M segmentations. Following [24], we evaluate SAM-Med2D both with and without adapter layers.
MedSAM [84]: MedSAM is a SAM ViT-B model fine-tuned with bounding box prompts on a collection of biomedical imaging datasets containing 1.5M image segmentation pairs. We evaluate MedSAM with bounding boxes only, because we found it to perform poorly when given point or mask prompts (Fig. 24).
@@
MIDeepSeg [82]: MIDeepSeg is an interactive segmentation framework designed to generalize to unseen tasks on medical images. MIDeepSeg takes interior margin points (positive clicks) as initial inputs. It crops the image based on those points and uses a CNN to make an initial prediction. Given additional clicks, the prediction can be refined using conditional random fields. We evaluate the pre-trained model, which was developed on placenta T2 MRI.
5 Evaluation
@@
We evaluate all methods using both manual and simulated interactions, with a focus on scribbles. Simulated interactions enable us to test on a large volume of images and tasks. However, simulations do not always match user behavior. We address this by (1) collecting a diverse dataset of manual scribbles for evaluation, and (2) conducting an interactive user study.
Figure 5: Example predictions. SP = ScribblePrompt. Top: predictions after one step of manual scribbles. Bottom: predictions after five steps of simulated interactions (one center click followed by one center correction click per step). See the Appendix E,F for more examples.
@@ 5.1 Manual Scribbles
Setup. We evaluate on two datasets of manual scribbles. First, we collected a dataset (MedScribble) of manual scribbles from three annotators for seven segmentation tasks from unseen datasets, with a total of 31 examples. The annotators were shown five training examples with the ground truth segmentation per task and instructed to draw positive and negative scribbles to indicate the region of interest on new images. We provide more details on MedScribble in Appendix E. Second, we used 380 slices from the ACDC dataset [14], which has scribble annotations for three labels and background.
In this evaluation, we make a prediction given a set of positive and negative scribbles, simulating a non-iterative scenario where the annotator draws several scribbles before running inference. We evaluate accuracy using Dice score [32] and 95th percentile Hausdorff Distance [52]. Since MIDeepSeg’s initial prediction network only takes positive inputs, we report results on both MIDeepseg’s initial prediction from the positive scribbles, and after applying MIDeepSeg’s refinement procedure using the negative scribbles as corrections. For SAM and variants fine-tuning SAM on clicks, we convert each scribble to a series of points. For MedSAM, we fit a bounding box to the positive scribbles.
Results. ScribblePrompt-UNet and ScribblePrompt-SAM produce the most accurate segmentations in a single step of manual scribbles on both our manual scribbles dataset and the ACDC scribbles dataset (Tab. 1). We show example predictions in Fig. 5 and Appendix E.
SAM and SAM-Med2D do not generalize well to scribble inputs (which they were not trained for). MedSAM has better predictions than other SAM baselines using the SAM architecture, however it is not able to make use of the negative scribbles and thus often misses segmentations with holes in them (Fig. 5). The initial predictions from MIDeepSeg’s network are poor, but improve after applying the refinement procedure.
Table 1: Manual scribbles. We report mean Dice and HD95 with 95% CI of segmentations predicted from manually-collected scribbles. Best and second best are highlighted.
MedScribble (n=93) ACDC (n=2,130)
Model $\uparrow$ Dice Score $\downarrow$ HD95   $\uparrow$ Dice Score $\downarrow$ HD95
SAM (ViT-b) $0.40\pm 0.05$ $32.58\pm 3.80$ $0.20\pm 0.01$ $108.39\pm 0.85$
SAM (ViT-h) $0.56\pm 0.05$ $14.61\pm 3.18$ $0.42\pm 0.02$ $59.78\pm 2.02$
SAM-Med2D w.o. adapter $0.52\pm 0.05$ $30.34\pm 3.29$ $0.16\pm 0.01$ $107.25\pm 1.12$
SAM-Med2D w/ adapter $0.30\pm 0.06$ $19.04\pm 3.63$ $0.17\pm 0.02$ $41.01\pm 3.04$
MIDeepSeg $0.68\pm 0.04$ $6.94\pm 1.06$ $0.58\pm 0.01$ $5.78\pm 0.26$
MIDeepSeg w/ refinement $0.81\pm 0.03$ $3.10\pm 0.67$ $0.73\pm 0.01$ $2.71\pm 0.18$
MedSAM (box) $0.70\pm 0.04$ $7.54\pm 1.35$ $0.70\pm 0.02$ $3.53\pm 0.28$
ScribblePrompt-SAM $0.87\pm 0.02$ $2.61\pm 0.84$ $0.77\pm 0.01$ $4.17\pm 0.32$
ScribblePrompt-UNet $0.84\pm 0.02$ $2.92\pm 0.88$ $0.84\pm 0.01$ $1.80\pm 0.11$
@@ 5.2 Simulated Interactions
Setup. We use our interaction simulator to evaluate the performance of ScribblePrompt and baselines with iterative scribbles across the 12 evaluation sets.
@@
We limit each scribble to cover a maximum of $w$ pixels, where $w$ is the maximum dimension of the image. For centerline and contour scribbles, a scribble might contain multiple disconnected components.
@@
  • Random Warm Start: Three random positive clicks and three random negative clicks to start, followed by one (positive or negative) correction click per step in the center of the largest component of the error region.
For each example and interaction procedure, we simulate a series of iterative interactions with five random seeds. Since MIDeepSeg cannot make accurate predictions with only a few clicks, we exclude it from simulations that start with a single click. For MedSAM, we show results for a bounding box prompt.
Results. Fig. 6 shows that both versions of ScribblePrompt outperform baseline methods for all simulated interaction procedures at all numbers of interactions. We show examples with simulated interactions in Fig. 1 and Fig. 5 (bottom). We show results with similar trends using bounding boxes and by dataset with fully-supervised baselines in Appendix F.
Discussion. Comparing the simulated scribble and click results (Fig. 6) highlights the efficiency of scribble-based interactions. Both ScribblePrompt models reach an average Dice above $0.8$ in two scribble steps. It takes five steps of carefully-placed center clicks or eight random clicks to reach the same average Dice score with ScribblePrompt-SAM.
Although SAM-Med2D was trained on a large biomedical imaging collection, it does not perform as well as SAM at refining its predictions on held-out evaluation sets. MedSAM only uses bounding box prompts and is not able to refine predictions, limiting its performance and usability.
Figure 6: Simulated clicks and scribbles. We simulate interactions following three scribble protocols and three click protocols. We show more results, with similar trends, in Sec. F.2 (Fig. 21).
5.3 User Study
Table 2: User study. We report means and standard deviations for Dice score, HD95 and total time per segmentation, as well as the median (and max) number of times the participants ($n=16$) refreshed the prediction.
SAM (ViT-b) ScribblePrompt-UNet